1 How much power will a centrifugal pump need in order to deliver 250 m³/h of Sulphuric acid (S.G. 1.83) to a head of 46m. Assume the pump efficiency is 80%.

2 If the pump in question “a” was actually supposed to deliver 270m³/h, what impeller diameter should you fit if the current impeller diameter is 380mm?
Answer #1

Sol 1

Discharge = 250m3/h or 250/3600 = 0.0694 m3/s

specific gravity = 1.83

so Density = 1.83 * 1000 = 1830 Kg/m3

Head = 46 m

so Power required = density * head * g * Discharge/pump efficiency = 1830*46*9.81* 0.0694/0.8 = 71684.53 W or 96.0918 hp (1hp =746 W) or 71.6845 KW Ans

Sol 2

now this required the concept of model testing

for centrifugal pump

(Q1/D13N1)m1 = (Q2/D23N2)m2

where m1 and m2 are first pump and second pump

Since nothing is said about speed N I am assuming it be constant for both the cases

N1=N2

Q2 = 0.075 m3/s (270/3600 = 0.075)

so D2=(D13*Q2/Q1)1/3= (0.383*0.075/0.0694)1/3

D2 = 389.95 mm or 390 mm

So new impleller diameter will be 390mm